The topic is " Find the Difference". For example 5-3= and example 14-(-2)= My answers is 2 for the first problem and 12 for the second problem, is that right?
Found 2 solutions by jim_thompson5910, midwood_trail:
Answer by jim_thompson5910(35256)

5-3=2 ... correct

14-(-2)= 14 + 2 = 16


Note: the two negatives cancel to a positive (ie they undo each other)

Answer by midwood_trail(310)

The topic is " Find the Difference". For example 5-3= and example 14-(-2)= My answers is 2 for the first problem and 12 for the second problem, is that right?
================================
To "find the difference" means to subtract or take away.
For the first problem, your son is right because 5 - 3 = 2.
For the second question, your son is wrong.
The reason he is wrong is because there is a NEGATIVE in front of the parentheses. See it?
This is the question:
14 - (-2)....This is read: "Fourteen minus negative two."
Believe it or not, there is an INVISIBLE -1 in front of the parentheses.
As you know, negative TIMES negative = positive.
We really have this:
14 minus -1 times -2 = 14 + 2 = 16
The answer to question 2 is 16 not 12.
Did you follow?
Tell your son to review the rules of signed numbers. If he does not take time out to memorize the rules of signed numbers, algebra will not be pretty.